Determine the values of the power supplied by each of the sources in the circuit shown bellow.

Determine the values of the power supplied by each
Answer #1

Solution: Apply KVL to the individual loops. 12 =101, +4012 12-101 + 24-(1-12 +0.6) 40 = 0 60 = 501, - 4012 Solving (1) and (